OpenStudy (anonymous):

Simplify the raidcal. 7 sqrt 72

OpenStudy (anonymous):

9*7 21sqrt(7) Not trying them?

OpenStudy (anonymous):

or even 9*8 :P

OpenStudy (anonymous):

7sqrt(72) 7sqrt(9*4*2) 42sqrt(2)

OpenStudy (anonymous):

i got 21 sqrt 8

OpenStudy (muhammad_nauman_umair):

42aqrt(2)

OpenStudy (anonymous):

8 is the product of a square and 2 so can be simplified further

OpenStudy (muhammad_nauman_umair):

42sqrt(2)

OpenStudy (callisto):

\[7\sqrt{72}=7\sqrt{3 \times 3 \times 2 \times 2 \times 2}=7\sqrt{3^2}\sqrt{2^3} = 7(3)(2)\sqrt{2}=42\sqrt{2}\]

OpenStudy (anonymous):

42 sqrt 2